# Prediction-Powered Adaptive Shrinkage Estimation

### v1

@@ 4.2 The Across Problem Adaptive Shrinkage Stage
5 Theoretical Results
In (18), we proposed selecting $\hat{\omega}$ by optimizing an unbiased surrogate of true risk. In this section we justify this procedure theoretically. Our first result establishes that CURE approximates the true loss (whose expectation is the compound risk in (9)) uniformly in $\omega$ as we consider more and more problems.
@@ 5 Theoretical Results
A principal consequence of Proposition 5.1 is that PAS with the data-driven choice of $\hat{\omega}$ in (18) has asymptotically smaller Bayes MSE (defined in (11)) than any of the estimators in (15), i.e. it has smaller MSE than both baselines in (2) as well as the PPI and PT estimators.
@@
and so
$\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\boldsymbol{\hat{\theta}}^{\mathrm{PAS}}_{\hat{\omega}}\Big{)}\leq\min\left\{\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\boldsymbol{\tilde{Z}}^{f}\Big{)},\,\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\Big{(}\boldsymbol{\hat{\theta}}^{\mathrm{PT}}\Big{)}\right\}+o(1).$
@@
In addition to the assumptions of Proposition 5.1, further assume that $N_{j}=\infty$ and that there exist $n\in\mathbb{N}$, $\tilde{\sigma}^{2}>0$ such that $n_{j}=n$ and $\tilde{\sigma}^{2}_{j}=\tilde{\sigma}^{2}$ for all $j$ almost surely. Let $\beta^{2}:=\mathbb{E}_{\mathbb{P}_{\eta}}[(\tilde{Z}^{f}_{j}-\theta_{j})^{2}]$ (which does not depend on $j$ as we are integrating over $\mathbb{P}_{\eta})$. Then, as $m\to\infty$,
$\mathcal{B}^{\mathbb{P}_{\eta}}_{m}\Big{(}\boldsymbol{\hat{\theta}}^{\mathrm{PAS}}_{\hat{\omega}}\Big{)}\leq\frac{\tilde{\sigma}^{2}\beta^{2}}{\tilde{\sigma}^{2}+\beta^{2}}+o(1).$
To interpret the result, it is instructive to compare the asymptotic upper bound on the MSE of PAS with the MSE in the last line of Table 1, i.e., with $(\sigma_{\varepsilon}^{2}\sigma_{\theta\mid\phi}^{2})/(\sigma_{\varepsilon}^{2}+\sigma_{\theta\mid\phi}^{2})$. Observe that $\tilde{\sigma}^{2}$ plays the role of $\sigma_{\varepsilon}^{2}$ (as already anticipated in (16)) which is smaller than the variance of the classical estimator (due to power tuning). Meanwhile, $\beta^{2}$ plays the role of $\sigma^{2}_{\theta\mid\phi}$. If the baseline $\tilde{Z}^{f}_{j}$ that is the mean of the ML predictions on the unlabeled datasets is doing a good job of predicting $\theta_{j}$, then $\beta^{2}$ will be small, and so PAS may have MSE substantially smaller than that of PT. On the other hand, even if $\beta^{2}$ is large (that is, even if the ML model is very biased), PAS asymptotically still has MSE less or equal than $\tilde{\sigma}^{2}$, the MSE of PT. We emphasize that the role of Proposition 5.3 is to provide intuition at the expense of strong assumptions. By contrast, the statement of Theorem 5.2 does not restrict heterogeneity (e.g., heteroscedasticity) across problems and allows for varying, finite unlabeled and labeled sample sizes.
@@ 6 Experiments
    prediction mean on unlabeled data;
@@
    the “shrink-average” estimator that shrinks $\hat{\theta}^{\mathrm{PT}}_{j}$ toward the PT group mean across all problems, $\bar{\theta}^{\mathrm{PT}}:=m^{-1}\sum_{j}\hat{\theta}^{\mathrm{PT}}_{j}$.
See Appendix C for detailed formulations and implementations for (v.) & (vi.) (the latter of which is inspired by the SURE-grand mean estimator of Xie et al. (2012)). The comparisons with (iv.) and (v.) also directly serve as ablation studies of the two stages in constructing the PAS estimator.
@@
We report the mean squared error (MSE) ($\pm$ 1 standard error) of each estimator $\boldsymbol{\hat{\theta}}$ by averaging $\tfrac{1}{m}\sum_{j=1}^{m}(\hat{\theta}_{j}-\theta_{j})^{2}$ across $K=200$ Monte Carlo replicates. In the synthetic model, we sample $\eta_{j}$ (and thus $\theta_{j}$) from the known prior $\mathbb{P}_{\eta}$. For the real-world datasets, since $\mathbb{P}_{\eta}$ is unknown, we follow the standard evaluation strategy in the PPI literature: we start with a large labeled dataset and use it to compute a pseudo-ground truth for each mean $\theta_{j}$. Then in each Monte Carlo replicate, we randomly split the data points of each problem into labeled/unlabeled partitions (where we choose a 20/80 split ratio). We provide more details on the benchmarking procedure for the MSE in Appendix D.
For real-world datasets, we introduce a second metric to assess whether improvements in MSE are driven by a few difficult problems rather than consistent performance gains: the percentage of problems improved relative to the classical estimator. This metric is defined as
%$\%\>\ \mathrm{Improved}\uparrow\;\;\,:=\;\;\,\frac{1}{m}\sum_{j=1}^{m}\mathbf{1}\left[(\hat{\theta}_{j}-\theta_{j})^{2}<(\hat{\theta}_{j}^{\mathrm{Classical}}-\theta_{j})^{2}\right]\times(100\%).$
Larger values of this metric are preferable.
@@ 6.1 Synthetic Model
Table 2: MSE ($\pm$ standard error) of different estimators under the synthetic model with predictors $f_{1}(x)=x^{2}$, respectively $f_{2}(x)=|x|$.
@@
For both predictors, we see that PAS outperforms all the baselines. With a good predictor $f_{1}$, both the prediction mean and the shrinkage estimator closely track PAS; in contrast, the PPI and PT estimators fail to fully leverage the accurate predictions, as their design enforces unbiasedness. The situation reverses for the less reliable predictor $f_{2}$: the prediction mean and the shrinkage estimator have high MSE, while estimators with built-in de-biasing mechanisms demonstrate greater resilience. PAS adapts effectively across these extremes, making it a robust choice for a wide range of problems and predictors.
@@ 6.2 Real-World Datasets
Table 3: Results aggregated over $K=200$ replicates on the Amazon review dataset with BERT-base and BERT-tuned predictors. Metrics are reported with $\pm$ 1 standard error.
Amazon (base $f$) Amazon (tuned $f$)
Estimator MSE (×$(\times 10^{-3})$ Improved $\uparrow$ MSE (×$(\times 10^{-3})$ Improved $\uparrow$
Classical 24.305 $\pm$ 0.189 baseline 24.305 $\pm$ 0.189 baseline
Prediction Avg 41.332 $\pm$ 0.050 30.7 $\pm$ 0.2 3.945 $\pm$ 0.011 75.4 $\pm$ 0.2
PPI 11.063 $\pm$ 0.085 62.4 $\pm$ 0.2 7.565 $\pm$ 0.066 70.4 $\pm$ 0.2
PT 10.633 $\pm$ 0.089 70.3 $\pm$ 0.2 6.289 $\pm$ 0.050 76.0 $\pm$ 0.2
Shrink Classical 15.995 $\pm$ 0.121 56.4 $\pm$ 0.3 3.828 $\pm$ 0.039 78.9 $\pm$ 0.2
Shrink Avg 9.276 $\pm$ 0.078 70.4 $\pm$ 0.2 6.280 $\pm$ 0.058 77.1 $\pm$ 0.2
PAS (ours) 8.517 $\pm$ 0.071 71.4 $\pm$ 0.2 3.287 $\pm$ 0.024 80.8 $\pm$ 0.2
Table 4: Results aggregated over $K=200$ replicates on the Galaxy dataset with ResNet50 predictor. Metrics are reported with $\pm$ 1 standard error.
Galaxy
Estimator MSE (×$(\times 10^{-3})$ Improved $\uparrow$
Classical 2.073 $\pm$ 0.028 baseline
Prediction Avg 7.195 $\pm$ 0.008 17.0 $\pm$ 0.2
PPI 1.149 $\pm$ 0.017 59.4 $\pm$ 0.3
PT 1.026 $\pm$ 0.015 67.7 $\pm$ 0.3
Shrink Classical 1.522 $\pm$ 0.016 48.8 $\pm$ 0.4
Shrink Avg 0.976 $\pm$ 0.014 68.9 $\pm$ 0.3
PAS (ours) 0.893 $\pm$ 0.011 67.3 $\pm$ 0.4
@@
We next evaluate PAS on two large-scale real-world datasets, highlighting its ability to leverage state-of-the-art deep learning models in different settings. We include only the essential setup below and defer additional data and model details (e.g., hyper-parameters, preprocessing) to Appendix D.
Amazon Review Ratings (SNAP, 2014).
@@
Many commercial and scientific studies involve collecting a large corpus of text and estimating an average score (rating, polarity, etc.) from it. A practitioner would often combine limited human annotations with massive automatic evaluations from ML models (Tyser et al., 2024; Baly et al., 2020). To emulate this setup, we consider mean rating estimation problems using the Amazon Fine Food Review dataset from Kaggle, where we artificially hide the labels in a random subset of the full data to serve as the unlabeled partition. Concretely, we estimate the average rating for the top $m=200$ products with the most reviews (from $\sim$200 to $\sim$900). For the $i$-th review of the $j$-th product, the covariate $X_{ij}$ consists of the review’s title and text concatenated, while the outcome $Y_{ij}$ is the star rating in $\{1,\dots,5\}$. We employ two black-box predictors: (1) BERT-base, a language model without fine-tuning (Devlin, 2018) and (2) BERT-tuned which is the same model but fine-tuned on a held-out set of reviews from other products. Neither of these models have seen the reviews for the 200 products during training.
Spiral Galaxy Fractions (Willett et al., 2013).
@@
The Galaxy Zoo 2 project contains the classification results of galaxy images from the the Sloan Digital Sky Survey (York et al., 2000, SDSS). We are interested in estimating the fraction of galaxies that are classified as “spiral,” i.e., have at least one spiral arm. The covariates $X_{ij}$ in this applications are images (we provide some examples in Figure 6 of the appendix). Existing PPI papers have focused on estimating the overall fraction; we demonstrate how this dataset’s metadata structure enables compound estimation of spiral fractions across distinct galaxy subgroups. We first use a pre-defined partition of the galaxies into 122 subgroups that is based on the metadata attribute REDSHIFT. Second, we estimate the fraction of spiral galaxies in all of the galaxy subgroups simultaneously. For the predictor, we train a ResNet50 convolutional neural network on a held-out set with around 50k images.
@@
For both datasets, we randomly split the data for each problem (a food product or galaxy subgroup) into a labeled and unlabeled partition with a 20/80 ratio. We repeat the random splitting $K=200$ times and report metrics averaged over all splits. Results are displayed in Table 3, and here is a brief summary:
@@
  • Amazon Review: similar to the trend in the synthetic model, the more accurate BERT-tuned model enables stronger shrinkage for PAS while the biased BERT-base predictions necessitate less shrinkage. Our PAS estimator adapts to both predictors and outperforms other baselines. PAS has the lowest MSE without sacrificing performance on our second metric.
@@
  • Galaxy Zoo 2: the predictions from ResNet50 are suboptimal, so the variance-reduction from power tuning dominates any benefit from shrinkage. PAS achieves the lowest MSE among all estimators, and improves individual estimates at a level on par with the shrink-towards-mean estimator.
In Figure 4, we visualize the PAS and PT estimators against the pseudo-ground truth $\theta_{j}$ for each problem in the Galaxy Zoo and Amazon Review (with BERT-tuned) datasets. With the quantitative results obtained above, it is not surprising to see PAS estimators being closer to the $45^{\circ}$ line of perfect estimation. More importantly, the spread of pseudo-ground truth $\theta_{j}$ suggests that there is substantial heterogeneity (in outcome means) across different problems, justifying the stratification procedure and the compound mean estimation setting.
Figure 4: Comparison between the PAS and PT (PPI++) estimators for each problem in the Galaxy Zoo and Amazon Review (with BERT-tuned) datasets. The dashed $45^{\circ}$ line indicates perfect estimation against the pseudo-ground truth $\theta_{j}$.

### v2

@@ 5 Asymptotic Optimality
A principal consequence of Proposition 5.1 is that PAS with the data-driven choice of $\hat{\omega}$ in (18) has asymptotically smaller Bayes MSE (defined in (11)) than any of the estimators in (15), i.e., it has smaller MSE than both baselines in (2) as well as the PPI and PT estimators.
@@
so $\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\big{(}\boldsymbol{\hat{\theta}}^{\mathrm{PAS}}_{\hat{\omega}}\big{)}\leq\min\big{\{}\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\big{(}\boldsymbol{\tilde{Z}}^{f}\big{)},\,\mathcal{B}_{m}^{\mathbb{P}_{\eta}}\big{(}\boldsymbol{\hat{\theta}}^{\mathrm{PT}}\big{)}\big{\}}+o(1).$
@@
In addition to the assumptions of Proposition 5.1, further assume that $N_{j}=\infty$ and that there exist $n\in\mathbb{N}$, $\tilde{\sigma}^{2}>0$ such that $n_{j}=n$ and $\tilde{\sigma}^{2}_{j}=\tilde{\sigma}^{2}$ for all $j$ almost surely. Let $\beta^{2}:=\mathbb{E}_{\mathbb{P}_{\eta}}[(\tilde{Z}^{f}_{j}-\theta_{j})^{2}]$ (which does not depend on $j$ as we are integrating over $\mathbb{P}_{\eta})$. Then,
$\mathcal{B}^{\mathbb{P}_{\eta}}_{m}\Big{(}\boldsymbol{\hat{\theta}}^{\mathrm{PAS}}_{\hat{\omega}}\Big{)}\leq\frac{\tilde{\sigma}^{2}\beta^{2}}{\tilde{\sigma}^{2}+\beta^{2}}+o(1)\,\text{ as }\,m\to\infty.$
To interpret the result, it is instructive to compare the asymptotic upper bound on the MSE of PAS with the MSE in the last line of Table 1, i.e., with $(\sigma_{\varepsilon}^{2}\sigma_{\theta\mid\phi}^{2})/(\sigma_{\varepsilon}^{2}+\sigma_{\theta\mid\phi}^{2})$. Observe that $\tilde{\sigma}^{2}$ plays the role of $\sigma_{\varepsilon}^{2}$ (as already anticipated in (16)) which is smaller than the variance of the classical estimator (due to power tuning). Meanwhile, $\beta^{2}$ plays the role of $\sigma^{2}_{\theta\mid\phi}$. If the baseline $\tilde{Z}^{f}_{j}$ (that is, the mean of the ML predictions on the unlabeled datasets) is doing a good job of predicting $\theta_{j}$, then $\beta^{2}$ will be small, and so PAS may have MSE substantially smaller than that of PT. On the other hand, even if $\beta^{2}$ is large (that is, even if the ML model is very biased), PAS asymptotically still has MSE less than or equal to $\tilde{\sigma}^{2}$, the MSE of PT. We emphasize that the role of Proposition 5.3 is to provide intuition at the expense of strong assumptions. [^12] We further elaborate on this proposition in Section A.2.3. By contrast, the statement of Theorem 5.2 does not restrict heterogeneity (e.g., heteroscedasticity) across problems and allows for varying, finite unlabeled and labeled sample sizes.
6 PAS with Unknown Second Moments
So far we have assumed that the second moments $\sigma_{j}^{2},\tau_{j}^{2},\gamma_{j}$ in (7) are known. In practice, e.g., in our numerical experiments with real-world datasets, we use sample-based estimates $\hat{\sigma}_{j}^{2},\hat{\tau}_{j}^{2}$ and $\hat{\gamma}_{j}$ instead (Section C.1). This approach works well empirically, even for small $n_{j}$, but is not covered by the theory above. To address this shortcoming, we develop variants UniPT (Section C.2) and UniPAS (Section C.3) that provide asymptotic guarantees (as $m\to\infty$) without requiring known or consistently estimable second moments ($n_{j},N_{j}$ remain bounded). Briefly, UniPT targets an optimal single power-tuning parameter $\lambda$ across all problems, and does so consistently as $m\to\infty$. UniPAS then builds upon UniPT, applying adaptive shrinkage similar to PAS, but with mechanisms to handle the unknown moments. These extensions are justified by theoretical results akin to those for PAS (cf. Theorem 5.2). In our experiments below, UniPAS is competitive with PAS.
@@ 7 Experiments
    the prediction mean on unlabeled data;
@@
    the “shrink-average” estimator that shrinks $\hat{\theta}^{\text{PT}}_{j}$ toward the PT group mean across all problems, i.e., toward $\bar{\theta}^{\mathrm{PT}}:=\tfrac{1}{m}\sum_{j=1}^{m}\hat{\theta}^{\text{PT}}_{j}$.
  7. (vii.)
    the UniPT and UniPAS estimators introduced in Section 6 and detailed in Appendix C.
We include the estimators in (vii.) only for the real-world experiments, as they are specifically designed for settings where the second moments are unknown. See Appendix D for detailed formulations and implementations for (v.) & (vi.) (the latter of which is inspired by the SURE-grand mean estimator of Xie et al. (2012)). The comparisons with (iv.) and (v.) also directly serve as ablation studies of the two stages in constructing the PAS estimator.
@@
We report the mean squared error (MSE) ($\pm$ 1 standard error) of each estimator $\boldsymbol{\hat{\theta}}$ by averaging $\tfrac{1}{m}\sum_{j=1}^{m}(\hat{\theta}_{j}-\theta_{j})^{2}$ across $K=200$ Monte Carlo replicates. In the synthetic model, we sample $\eta_{j}$ (and thus $\theta_{j}$) from the known prior $\mathbb{P}_{\eta}$. For the real-world datasets, since $\mathbb{P}_{\eta}$ is unknown, we follow the standard evaluation strategy in the PPI literature: we start with a large labeled dataset and use it to compute a pseudo-ground truth for each mean $\theta_{j}$. Then in each Monte Carlo replicate, we randomly split the data points of each problem into labeled/unlabeled partitions (where we choose a 20/80 split ratio). [^13] In Appendix E, we vary this ratio from 1% to 40%, and provide more details on the benchmarking procedure for MSE.
For real-world datasets, we introduce a second metric to assess whether improvements in MSE are driven by a few difficult problems rather than consistent performance gains: the percentage of problems improved relative to the classical estimator (abbreviated as “$\mathrm{Improved}\,\%\,\uparrow$”). This metric is defined as $\frac{1}{m}\sum_{j=1}^{m}\mathbf{1}[(\hat{\theta}_{j}-\theta_{j})^{2}<(\hat{\theta}_{j}^{\mathrm{Classical}}-\theta_{j})^{2}]\times(100\%).$ Larger values of this metric are preferable.
@@ 7.1 Synthetic Model
Table 2: MSE ($\pm$ standard error) of different estimators under the synthetic model with predictors $f_{1}(x)=x^{2}$ and $f_{2}(x)=|x|$.
@@
For both predictors, we see that PAS outperforms all the baselines. With a good predictor $f_{1}$, both the prediction mean and the shrinkage estimator closely track PAS; in contrast, the PPI and PT estimators fail to fully leverage the accurate predictions, as their design enforces unbiasedness. The situation reverses for the less reliable predictor $f_{2}$: the prediction mean and the shrinkage estimator have high MSE, while estimators with built-in debiasing mechanisms demonstrate greater resilience. PAS adapts effectively across these extremes, making it a handy choice for a wide range of problems and predictors.
Table 3: Results aggregated over $K=200$ replicates on three real-world datasets: Amazon review ratings with BERT-base and BERT-tuned predictors, and spiral galaxy fractions with ResNet50 predictor. The bottom two rows are the UniPT and UniPAS estimators detailed in Appendix C as variants of PT and PAS. Metrics are reported with $\pm$ 1 standard error.
Amazon (base $f$) Amazon (tuned $f$) Galaxy Estimator MSE (×$(\times 10^{-3})$ Improved $\uparrow$ MSE (×$(\times 10^{-3})$ Improved $\uparrow$ MSE (×$(\times 10^{-3})$ Improved $\uparrow$ Classical 24.305 $\pm$ 0.189 baseline 24.305 $\pm$ 0.189 baseline 2.073 $\pm$ 0.028 baseline Prediction Avg 41.332 $\pm$ 0.050 30.7 $\pm$ 0.2 3.945 $\pm$ 0.011 75.4 $\pm$ 0.2 7.195 $\pm$ 0.008 17.0 $\pm$ 0.2 PPI 11.063 $\pm$ 0.085 62.4 $\pm$ 0.2 7.565 $\pm$ 0.066 70.4 $\pm$ 0.2 1.149 $\pm$ 0.017 59.4 $\pm$ 0.3 PT 10.633 $\pm$ 0.089 70.3 $\pm$ 0.2 6.289 $\pm$ 0.050 76.0 $\pm$ 0.2 1.026 $\pm$ 0.015 67.7 $\pm$ 0.3 Shrink Classical 15.995 $\pm$ 0.121 56.4 $\pm$ 0.3 3.828 $\pm$ 0.039 78.9 $\pm$ 0.2 1.522 $\pm$ 0.016 48.8 $\pm$ 0.4 Shrink Avg 9.276 $\pm$ 0.078 70.4 $\pm$ 0.2 6.280 $\pm$ 0.058 77.1 $\pm$ 0.2 0.976 $\pm$ 0.014 68.9 $\pm$ 0.3 PAS (ours) 8.517 $\pm$ 0.071 71.4 $\pm$ 0.2 3.287 $\pm$ 0.024 80.8 $\pm$ 0.2 0.893 $\pm$ 0.011 67.3 $\pm$ 0.4 UniPT (ours) 10.272 $\pm$ 0.084 70.0 $\pm$ 0.2 6.489 $\pm$ 0.053 76.2 $\pm$ 0.2 1.017 $\pm$ 0.015 67.7 $\pm$ 0.3 UniPAS (ours) 8.879 $\pm$ 0.073 69.5 $\pm$ 0.2 3.356 $\pm$ 0.031 77.6 $\pm$ 0.2 0.909 $\pm$ 0.011 66.7 $\pm$ 0.3
@@ 7.2 Real-World Datasets
We next evaluate PAS on two large-scale real-world datasets, highlighting its ability to leverage state-of-the-art deep learning models in different settings. [^14] We include only the essential setup below and defer additional data and model details (e.g., hyper-parameters, preprocessing) to Appendix E.
@@
Amazon Review Ratings (SNAP, 2014).   Many commercial and scientific studies involve collecting a large corpus of text and estimating an average score (rating, polarity, etc.) from it. A practitioner would often combine limited expert annotations with massive automatic evaluations from ML models (Baly et al., 2020; Egami et al., 2023; Fan et al., 2024; Mozer & Miratrix, 2025). To emulate this setup, we consider mean rating estimation problems using the Amazon Fine Food Review dataset from Kaggle, where we artificially hide the labels in a random subset of the full data to serve as the unlabeled partition. Concretely, we estimate the average rating for the top $m=200$ products with the most reviews (from $\sim$200 to $\sim$900). For the $i$-th review of the $j$-th product, the covariate $X_{ij}$ consists of the review’s title and text concatenated, while the outcome $Y_{ij}$ is the star rating in $\{1,\dots,5\}$. We employ two black-box predictors: (1) BERT-base, a language model without fine-tuning (Devlin et al., 2019) and (2) BERT-tuned which is the same model but fine-tuned on a held-out set of reviews from other products. Neither of them are trained on the reviews of the 200 products we evaluate.
@@
Spiral Galaxy Fractions (Willett et al., 2013).   The Galaxy Zoo 2 project contains the classification results of galaxy images from the the Sloan Digital Sky Survey (York et al., 2000, SDSS). We are interested in estimating the fraction of galaxies that are classified as “spiral,” i.e., have at least one spiral arm. The covariates $X_{ij}$ in this applications are images (we provide some examples in Figure 5 of the appendix). Existing PPI papers have focused on estimating the overall fraction; we demonstrate how this dataset’s metadata structure enables compound estimation of spiral fractions across distinct galaxy subgroups. We first use a pre-defined partition of the galaxies into 100 subgroups that is based on the metadata attribute WVT_BIN. Second, we estimate the fraction of spiral galaxies in all of the galaxy subgroups simultaneously. The predictor is a ResNet50 network trained on a held-out set with around 50k images.
@@
For both datasets, we randomly split the data for each problem (a food product or galaxy subgroup) into a labeled and unlabeled partition with a 20/80 ratio. We repeat the random splitting $K=200$ times and report metrics averaged over all splits. Here we summarize the results in Table 3:
@@
Amazon Review: similar to the trend in the synthetic model, the more accurate BERT-tuned model enables stronger shrinkage for PAS while the biased BERT-base predictions necessitate less shrinkage. Our PAS estimator adapts to both predictors and outperforms other baselines. PAS has the lowest MSE and highest % Improved $\uparrow$.
@@
Galaxy Zoo 2: the predictions from ResNet50 are suboptimal, so the variance-reduction from power tuning dominates any benefit from shrinkage. PAS achieves the lowest MSE among all estimators, and improves individual estimates at a level on par with the shrink-average estimator.